# Flows of linear orders on sparse graphs

### v1

@@ 1. Introduction
The paper [13] of Kechris, Pestov and Todorčević established links between topological dynamics and structural Ramsey theory, with further developments in [17], [19], [2] (among others). We assume the reader is familiar with the background here, and briefly recall three key results, which we formulate for strong classes (classes of structures where we restrict to a particular subclass of permitted embeddings – see [4, Definition 2.1]):
@@
Theorem ([13, Theorem 10.8], [17, Theorem 5], [19, Theorem 5.7]).
@@
The paper [4], which was the starting point for the current paper, showed that classes of sparse graphs used in Hrushovski constructions ([10], [11]) demonstrate different behaviour to classes previously studied in the KPT context. A graph $A$ is $k$-sparse if for all finite $B\subseteq A$, we have $|E(B)|\leq k|B|$. It is well-knwon ([15]) that a graph is $k$-sparse iff it is $k$-orientable. We take $k=2$ for presentational simplicity.
We briefly describe the classes $\mathcal{C}_{0},\mathcal{C}_{1},\mathcal{C}_{F}$ of sparse graphs found in [4]. Let $\mathcal{C}_{0}$ denote the class of finite $2$-sparse graphs. For $A,B\in\mathcal{C}_{0}$, we write $A\leq_{s}B$ if there exists a $2$-orientation of $B$ in which $A$ is successor-closed (by [6, Lemma 1.5], this is equivalent to another phrasing in terms of predimension). With this notion of $\leq_{s}$-substructure, we have the free amalgamation class $(\mathcal{C}_{0},\leq_{s})$ with Fraïssé limit $M_{0}$ (this structure, an “ab initio Hrushovski construction”, was first described in [11]).
We also have a “simplified” version of $M_{0}$, first considered in [6]. Let $\mathcal{D}_{1}$ denote the class of finite $2$-oriented graphs with no directed cycles, and let $\mathcal{C}_{1}$ be the class of graph reducts of structures in $\mathcal{D}_{1}$. For $A,B\in\mathcal{C}_{1}$, write $A\leq_{1}B$ if there exists an expansion $B^{+}\in\mathcal{D}_{1}$ in which $A$ is successor-closed. Then $(\mathcal{C}_{1},\leq_{1})$ is again a free amalgamation class, whose Fraïssé limit we denote by $M_{1}$.
@@
Equivalently, by [13, Theorem 4.8], we have that $M$ has no coprecompact Ramsey expansion.
@@
Note that if $M(G)$ has a comeagre orbit, then so does any minimal $G$-flow (see [1]), so the above result shows again that $\operatorname{Aut}(M)$ for $M=M_{1},M_{0},M_{F}$ has non-metrisable universal minimal flow, using [2, Theorem 1.2]. In the context of the above result, T. Tsankov asked the following ([4, concluding remarks]):
@@
David Evans suggested that the author investigate the $\operatorname{Aut}(M)$-flow $\mathcal{LO}(M)$ of linear orders on $M$. We obtain the following result (the main result of this paper), for $M_{1}$, the “simplified version” of $M_{0}$ (see 2.4):
@@
This result demonstrates that the phenomenon seen in [4, Theorem 5.2] occurs more generally for other flows on $M_{1}$, partially answering the question of Tsankov. We also note in passing that $M_{1}$ is $\omega$-saturated and its theory is $\omega$-stable (see [6]).
To prove 4.1, we take the class of finite ordered graphs which $\leq$-embed into some element of the minimal flow $Y$, and show that this class fails to have the weak amalgamation property – this gives the result, using 2.32. To show failure of the weak amalgamation property, we will use the Ramsey expansion given by the admissible orders, from [5, Section 3.1]. We discuss these in Section 3.
@@
Recall that we consider $M_{1}$ to be a “simplified version” of $M_{0}$. It would be interesting to know if there is an analogous “simplified version” of $M_{F}$ – if so, it may be possible to prove 4.1 for an $\omega$-categorical structure. See [18, Chapter 7].
@@ 2. Background
We briefly summarise the background required for Section 3 and 4.1. This section is mostly based on [4, Section 2].
@@
2.1. Graphs and oriented graphs: notation and setup
A graph $(A,\operatorname{E}(A))$ consists of a vertex set $A$ and an edge set $\operatorname{E}(A)\subseteq A^{(2)}$ (where $A^{(2)}$ denotes the set of unordered pairs). Here we only work with simple graphs: graphs with no loops or multiple edges. A set $\rho_{A}\subseteq A^{2}$ is an orientation of $(A,\operatorname{E}(A))$ if:
  • for $xy\in\operatorname{E}(A)$, exactly one of $(x,y),(y,x)$ is in $\rho_{A}$;
  • for $(x,y)\in\rho_{A}$, $xy\in\operatorname{E}(A)$.
We refer to $(A,\operatorname{E}(A),\rho_{A})$ as an oriented graph. We usually just write $(A,\rho_{A})$ or $A$, the edge set being clear from context.
We work with graphs in first-order languages as usual: the language of graphs $L_{\text{graph}}=\{E\}$ with $E$ binary, and a $L_{\text{graph}}$-structure $(A,E_{A})$ is then a graph if $E_{A}\subseteq A^{2}$ is symmetric and irreflexive. We expand $L_{\text{graph}}$ to the language of oriented graphs $L_{\text{or}}=\{E,\rho\}$, where $\rho$ is binary. A $L_{\text{or}}$-structure $(A,E_{A},\rho_{A})$ is an oriented graph if $\rho_{A}$ is an orientation of the graph $(A,E(A))$, where $E(A)$ is the edge set of unordered pairs induced by $E_{A}$. When we refer to a subgraph of a graph, or an oriented subgraph of an oriented graph, we mean an $L_{\text{graph}}$- or $L_{\text{or}}$-substructure respectively.
Definition 2.1.
Let $(A,\rho_{A})$ be an oriented graph. If $(x,y)\in\rho_{A}$, we write $xy\in\rho_{A}$. We call $xy$ an out-edge of $x$ and an in-edge of $y$. We call $y$ an out-vertex or successor of $x$, and $x$ an in-vertex or predecessor of $y$. The out-neighbourhood $\operatorname{N_{{}_{\!+}}}(x)$ and in-neighbourhood $\operatorname{N_{{}_{\!--}}}(x)$ of $x$ consist of the out-vertices and in-vertices respectively of $x$, and we define the out-degree $\operatorname{d_{{}_{+}}\!}(x)=|\operatorname{N_{{}_{\!+}}}(x)|$ and the in-degree $\operatorname{d_{{}_{--}}\!}(x)=|\operatorname{N_{{}_{\!--}}}(x)|$.
@@ 2.2. Sparse graphs: $\mathcal{C}_{0}$ and $\mathcal{C}_{1}$
Definition 2.2.
Let $k\in\mathbb{N}_{+}$. A graph $A$ is $k$-sparse if for all $B\subseteq_{\text{fin.\!}}A$, $|\operatorname{E}(B)|\leq k|B|$.
Let $(A,\rho_{A})$ be an oriented graph. We call $\rho_{A}$ a $k$-orientation if for $x\in A$, $|\operatorname{N_{{}_{\!+}}}(x)|\leq k$. If an undirected graph $A$ has a $k$-orientation, we say it is $k$-orientable.
The following well-known proposition ([15]) will be a key tool here.
Proposition 2.3 ([4], Th. 3.4).
A graph $A$ is $k$-orientable iff it is $k$-sparse.
For presentational simplicity, we will usually work with $k=2$. Our results generalise straightforwardly for $k>2$.
Definition 2.4.
Let $\mathcal{C}_{0}$ be the class of finite $2$-sparse graphs, and let $\mathcal{D}_{0}$ be the class of finite $2$-oriented graphs. By Proposition 2.3 we have that $\mathcal{C}_{0}$ is the class of graph reducts of $\mathcal{D}_{0}$.
Let $\mathcal{D}_{1}$ be the class of finite $2$-oriented graphs with no directed cycles. By a slight abuse of terminology, we will refer to a $2$-oriented graph with no directed cycles as an acyclic $2$-oriented graph. Let $\mathcal{C}_{1}$ be the class of graph reducts of $\mathcal{D}_{1}$.
We may also define $\mathcal{C}_{1}$ directly, as per the lemma below.
Lemma 2.5 ([6], Lem. 1.3).
A finite graph $A$ has a $k$-orientation with no directed cycles iff every non-empty subgraph $B$ has a vertex of degree $\leq k$ in $B$.
Proof.
:absent:$\Rightarrow$:⇒ : Consider the induced orientation on $B$ from an acyclic $k$-orientation on $A$. As this orientation is acyclic, $B$ contains a vertex $v$ which has no in-edges in $B$, and therefore the degree of $v$ in $B$ is equal to its out-degree in $B$, which is $\leq k$.
:absent:$\Leftarrow$:⇐ : We use induction on $|A|$. For $|A|=1$, the claim is trivial. For the inductive step, take a vertex $a\in A$ of degree $\leq k$. By assumption, we may give $A\setminus\{a\}$ an acyclic $k$-orientation. Then orient the edges of $a$ outwards from $a$ to produce an acyclic $k$-orientation of $A$. ∎
@@
Thus $\mathcal{C}_{1}$ consists of the finite graphs $A$ where every non-empty subgraph $B\subseteq A$ has a vertex of degree $\leq 2$.
2.3. Strong classes
To produce amalgamation classes from the classes of structures defined in the previous section, we will consider certain distinguished subclasses of embeddings. We assume the reader is familiar with Fraïssé theory for strong classes: see [3, Section 3] for more details.
Definition 2.6.
Let $\mathcal{K}$ be a class of finite $L$-structures closed under isomorphisms. Let $\mathcal{S}\subseteq\operatorname{Emb}(\mathcal{K})$ be a class of embeddings between structures in $\mathcal{K}$ which contains all isomorphisms and is closed under composition, and which also satisfies:
  1. ()($\ast$)( ∗ )
    if $f:A\to C$ is in $\mathcal{S}$ and $f(A)\subseteq B\subseteq C$ with $B\in\mathcal{K}$, then $f:A\to B$ is in $\mathcal{S}$.
Then we call $(\mathcal{K},\mathcal{S})$ a strong class, and call the elements of $\mathcal{S}$ strong embeddings.
If $A,B\in\mathcal{K}$, $A\subseteq B$ and the inclusion map $\iota:A\hookrightarrow B$ is in $\mathcal{S}$, then we write $A\leq B$ and say $A$ is a strong substructure of $B$. We have that $\leq$ is reflexive and transitive, and if $A\leq C$ and $A\subseteq B\subseteq C$ with $B\in\mathcal{K}$, then $A\leq B$. We will often write $(\mathcal{K},\leq)$ instead of $(\mathcal{K},\mathcal{S})$, and we will use different symbols resembling $\leq$ (e.g. $\leq_{1},\leq_{s},\sqsubseteq_{s}$) to indicate different classes of embeddings.
Let $(\mathcal{K},\leq)$ be a strong class. Then we have that for $f:A\to B$ in $\mathcal{S}$, if $X\leq A$, then $f(X)\leq B$.
We now define strong substructures of infinite structures. Let $(\mathcal{K},\leq)$ be a strong class. Let $A_{1}\leq A_{2}\leq\cdots$ be an increasing $\leq$-chain of structures in $\mathcal{K}$ with union $M$. Let $A\subseteq_{\text{fin.\!}}M$. Then we write $A\leq M$ to mean that there is some $A_{i}$ ($i\geq 1$) with $A\leq A_{i}$. By ()($\ast$)( ∗ ), this definition is independent of the particular $\leq$-chain we consider.
Let $g\in\operatorname{Aut}(M)$. It is easy to see that if $A\leq M$, then $gA\leq M$: that is, all $g\in\operatorname{Aut}(M)$ preserve $\leq$.
2.4. Fraïssé theory for strong classes
The classical Fraïssé theory directly generalises to strong classes. (See [3, Section 3]. For the classical Fraïssé theory, originally developed in [8], see [9].)
Given a strong class $(\mathcal{K},\leq)$, we define the joint embedding property (JEP), amalgamation property (AP) and free amalgamation as before, but we restrict to $\leq$-embeddings.
Definition 2.7.
Let $(\mathcal{K},\leq)$ be a strong class, and let $M$ be a union of an increasing $\leq$-chain of structures in $\mathcal{K}$. Let $\operatorname{Age}_{\leq}(M)$ denote the class of finite structures which $\leq$-embed into $M$.
$M$ has the $\leq$-extension property if for all $A,B\in\operatorname{Age}_{\leq}(M)$ and $\leq$-embeddings $f:A\to M,g:A\to B$, there exists a $\leq$-embedding $h:B\to M$ with $hg=f$.
$M$ is $\leq$-homogeneous if each isomorphism $f:A\to A^{\prime}$ between finite $A,A^{\prime}\leq M$ extends to an automorphism of $M$.
Definition 2.8.
Let $(\mathcal{K},\leq)$ be a strong class. We say that $(\mathcal{K},\leq)$ is an amalgamation class or Fraïssé class if $(\mathcal{K}\,\leq)$ contains countably many isomorphism types, and has the joint embedding and amalgamation properties.
Theorem 2.9 (Fraïssé-Hrushovski).
Let $(\mathcal{K},\leq)$ be an amalgamation class. Then there is a structure $M$ which is a union of an increasing $\leq$-chain, such that $M$ is $\leq$-homogeneous and $\operatorname{Age}_{\leq}(M)=\mathcal{K}$. Such an $M$ is unique up to isomorphism.
We call this structure the Fraïssé limit of $\mathcal{K}$.
2.5. Strong expansions
We will often have the situation where we have a strong class of $L$-structures together with a strong class of $L^{+}$-structures (with a potentially different notion of distinguished embedding), where $L^{+}\supseteq L$ is a relational language expanded from $L$.
Definition 2.10 ([4], Def. 2.9).
Let $(\mathcal{K},\leq)$ be a strong class of $L$-structures, and let $(\mathcal{K}^{+},\leq^{+})$ be a strong class of $L^{+}$-structures. We call $(\mathcal{K}^{+},\leq^{+})$ is a strong expansion of $(\mathcal{K},\leq)$ if:
  1. (1)
    $\mathcal{K}$ is the class of $L$-reducts of $\mathcal{K}^{+}$;
  2. (2)
    for $\leq^{+}$-strong $f:A^{+}\to B^{+}$, $f:A^{+}|_{L}\to B^{+}|_{L}$ is $\leq$-strong;
  3. (3)
    for $\leq$-strong $f:A\to B$ and $A^{+}\in\mathcal{K}^{+}$ an expansion of $A$, there exists an expansion $B^{+}\in\mathcal{K}^{+}$ of $B$ such that $f:A^{+}\to B^{+}$ is $\leq^{+}$-strong.
Lemma 2.11.
Let $(\mathcal{K}^{+},\leq^{+})$ be a strong expansion of the strong class $(\mathcal{K},\leq)$. If $(\mathcal{K}^{+},\leq^{+})$ is a (free) amalgamation class, then so is $(\mathcal{K},\leq)$. If $M^{+}$ is the Fraïssé limit of $(\mathcal{K}^{+},\leq^{+})$, then $M=M^{+}|_{L}$ is the Fraïssé limit of $(\mathcal{K},\leq)$, and $\operatorname{Aut}(M^{+})$ is a closed subgroup of $\operatorname{Aut}(M)$.
The proof is straightforward.
Definition 2.12.
A particular case of Definition 2.10 is the order expansion of a strong class. Let $(\mathcal{K},\leq)$ be a strong class of $L$-structures, and let $L^{\prec}=L\cup\{\prec\}$ with $\prec$ binary. Let $\mathcal{K}^{\prec}$ be the class of $L^{+}$-structures $(A,\prec_{A})$, where $A\in\mathcal{K}$ and $\prec_{A}$ is a linear order on $A$. For $(A,\prec_{A}),(B,\prec_{B})\in\mathcal{K}^{\prec}$, write $(A,\prec_{A})\leq(B,\prec_{B})$ if $A\leq B$ and $\prec_{A}=\prec_{B}|_{A}$. We then have that $(\mathcal{K}^{\prec},\leq)$ is a strong class and a strong expansion of $(\mathcal{K},\leq)$. We call $(\mathcal{K}^{\prec},\leq)$ the order expansion of the strong class $(\mathcal{K},\leq)$.
@@ 2.6. Sparse graphs: $\sqsubseteq_{s}$ and $\leq_{1}$
We will now describe the distinguished notions of embedding involved in defining particular strong classes for $\mathcal{C}_{1}$ and $\mathcal{D}_{1}$. (The material in this section is based on Section 3.4 of [4], and for $\leq_{1}$ was originally developed in [7].)
Definition 2.13.
@@
Let $A$ be an oriented graph. For $B\subseteq A$, we write $B\sqsubseteq_{s}A$ to mean that $B$ is successor-closed in $A$: for $b\in B$, if $ba\in\rho_{A}$ then $a\in B$. Note that intersections of successor-closed subsets are also successor-closed. For $B\subseteq A$, the successor-closure $\operatorname{scl}(B)$ is the smallest successor-closed subset of $A$ containing $B$.
Definition 2.14.
@@
Let $A,B\in\mathcal{C}_{1}$ with $A\subseteq B$. We write $A\leq_{1}B$ if there exists an acyclic $2$-orientation $B^{+}\in\mathcal{D}_{1}$ of $B$ in which the induced orientation $A^{+}\in\mathcal{D}_{1}$ on $A$ has $A^{+}\sqsubseteq_{s}B^{+}$.
Lemma 2.15.
Let $B\in\mathcal{C}_{1}$.
  1. (1)
    $A\leq_{1}B$, $X\subseteq B$Rightarrow $A\cap X\leq_{1}X$.
  2. (2)
    $A\leq_{1}C\leq_{1}B$Rightarrow $A\leq_{1}B$.
  3. (3)
    $A_{1},A_{2}\leq_{1}B$Rightarrow $A_{1}\cap A_{2}\leq_{1}B$.
The above three statements also hold for $B\in\mathcal{D}_{1}$ and $\sqsubseteq_{s}$.
Proof.
(1): immediate. (2): let $B^{+}$ be an acyclic $2$-orientation of $B$ in which $C$ is successor-closed. Create a new acyclic $2$-orientation $B^{++}$ on $B$ by taking $B^{+}$ and replacing the induced orientation on $C$ with an acyclic $2$-orientation on $C$ in which $A$ is successor-closed. Then $A$ is successor-closed in $B^{++}$. (3): follows from (1) and (2).
For $B\in\mathcal{D}_{1}$ and $\sqsubseteq_{s}$, the proof of the above three statements is straightforward. ∎
By the above lemma (and similar arguments to the proof of (2) in Lemma 2.15), we have that:
Lemma 2.16.
@@
$(\mathcal{C}_{1},\leq_{1})$ and $(\mathcal{D}_{1},\sqsubseteq_{s})$ are strong classes with free amalgamation, and $(\mathcal{D}_{1},\sqsubseteq_{s})$ is a strong expansion of $(\mathcal{C}_{1},\leq_{1})$.
Let $M_{1}$ be the Fraïssé limit of $(\mathcal{C}_{1},\leq_{1})$ and let $G_{1}=\operatorname{Aut}(M_{1})$.
2.7. Topological dynamics and Ramsey classes
A $G$-flow is a continuous action $G\curvearrowright X$ of a Hausdorff topological group $G$ on a nonempty compact Hausdorff topological space $X$. We will often simply write $X$ to refer to the $G$-flow $G\curvearrowright X$ when this is clear from context. Given $G\curvearrowright X$, the orbit closure $\overline{G\cdot x}$ of a point $x\in X$ is a $G$-invariant compact subset of $X$. In general, a nonempty compact $G$-invariant subset $Y\subseteq X$ defines a subflow by restricting the $G$-action to $Y$. A $G$-flow on $X$ is minimal if it contains no proper subflows. A $G$-flow is minimal iff every orbit is dense. By Zorn’s lemma, every $G$-flow contains a minimal subflow.
Let $X,Y$ be $G$-flows. A $G$-flow morphism $X\to Y$ is a continuous map $\pi:X\to Y$ such that $\pi(g\cdot x)=g\cdot\pi(x)$ (this property is called $G$-equivariance). Bijective $G$-flow morphisms are isomorphisms, as they are between compact Hausdorff spaces. If $Y$ is minimal, then any $G$-flow morphism $X\to Y$ is surjective, as the image is a subflow.
Definition 2.17.
Let $G$ be a Hausdorff topological group. If every $G$-flow has a $G$-fixed point, we call $G$ extremely amenable.
We now define the Ramsey property for strong classes. (This section is based on Section 2.2 of [4].)
Definition 2.18.
Let $(\mathcal{K},\leq)$ be a strong class. For $A,B\in\mathcal{K}$, we write $\binom{B}{A}=\{A^{\prime}\leq B:A^{\prime}\cong A\}$ for the set of $\leq$-copies of $A$ inside $B$. (Note that by the definition of a strong class, if $B\leq C\in\mathcal{K}$, then $\binom{B}{A}=\binom{C}{A}\cap\mathcal{P}(B)$.)
For $A,B\in\mathcal{K}$, $r\in\mathbb{N}_{+}$, an $r$-colouring of the set $\binom{B}{A}$ is a function $\chi:\binom{B}{A}\to\{1,\cdots,r\}$. We say that $\binom{B}{A}$ is monochromatic in the $r$-colouring $\chi$ if $\chi$ is constant on $\binom{B}{A}$.
For $A,B,C\in\mathcal{K}$, $r\in\mathbb{N}_{+}$, we write $C\rightarrow(B)^{A}_{r}$ if for every $r$-colouring $\chi$ of $\binom{C}{A}$, there exists $B^{\prime}\in\binom{C}{B}$ such that $\binom{B^{\prime}}{A}$ is monochromatic in $\chi$.
We say that $(\mathcal{K},\leq)$ has the Ramsey property if for $r\in\mathbb{N}_{+}$, $A,B\in\mathcal{K}$, there exists $C\in\mathcal{K}$ with $C\rightarrow(B)^{A}_{r}$.
The following is a well-known observation of Nešetřil ([16]) which shows a strong connection between Fraïssé theory and structural Ramsey theory (see [4, Theorem 2.11] for a proof):
Proposition 2.19.
Let $(\mathcal{K},\leq)$ be a strong, rigid class with JEP and the Ramsey property. Then $(\mathcal{K},\leq)$ has the amalgamation property.
The following is one of the key results of [13], and establishes a link between topological dynamics and structural Ramsey theory:
Theorem 2.20 ([13], Th. 4.8; [17], Th. 1).
Let $(\mathcal{K},\leq)$ be an amalgamation class of rigid $L$-structures, with Fraïssé limit $M$ and $G=\operatorname{Aut}(M)$. Then $G$ is extremely amenable iff $(\mathcal{K},\leq)$ has the Ramsey property.
Here, a finite $L$-structure is rigid if it has trivial automorphism group.
2.8. Reasonable expansions and the expansion property
In this section, $L^{+}$ is an expansion of the language $L$.
Here we follow Section 2.3 of [4], which takes the notion of a reasonable expansion from [19].
Definition 2.21.
Let $(\mathcal{K},\leq)$ be an amalgamation class of $L$-structures. A class $\mathcal{D}$ of finite $L^{+}$-structures is a reasonable expansion of $(\mathcal{K},\leq)$ if $\mathcal{D}$ is closed under isomorphisms and satisfies the following:
  1. (1)
    $\mathcal{K}$ is the class of $L$-reducts of $\mathcal{D}$;
  2. (2)
    for $A\in\mathcal{K}$, $A$ has finitely many expansions in $\mathcal{D}$ (weak coprecompactness);
  3. (3)
    for $B^{+}\in\mathcal{D}$, if $A^{+}\subseteq B^{+}$ and $A^{+}|_{L}\leq B^{+}|_{L}$, then $A^{+}\in\mathcal{D}$;
  4. (4)
    for $f:A\to B$ a strong embedding in $(\mathcal{K},\leq)$, if $A^{+}\in\mathcal{D}$ is an expansion of $A$, then there exists an expansion $B^{+}\in\mathcal{D}$ of $B$ such that $f:A^{+}\to B^{+}$ is an embedding.
Note that if $(\mathcal{K}^{+},\leq^{+})$ is a strong expansion of $(\mathcal{K},\leq)$, then $\mathcal{K}^{+}$ satisfies properties (1) and (4) above.
Lemma 2.22.
For any amalgamation class $(\mathcal{K},\leq)$ of $L$-structures, the order expansion $\mathcal{K}^{\prec}$ of $(\mathcal{K},\leq)$ (see Definition 2.12) is a reasonable expansion of the class $(\mathcal{K},\leq)$.
The proof is straightforward.
Lemma 2.23.
The class $\mathcal{D}_{1}$ is a reasonable expansion of $(\mathcal{C}_{1},\leq_{1})$.
Proof.
As $(\mathcal{D}_{1},\leq_{s})$ is a strong expansion of $(\mathcal{C}_{1},\leq_{1})$, we have that $\mathcal{D}_{1}$ satisfies (1) and (4) in the definition of reasonableness. Parts (2) and (3) are immediate. ∎
Definition 2.24.
Let $(\mathcal{K},\leq)$ be an amalgamation class with Fraïssé limit $M$. Let $\mathcal{D}$ be a reasonable $L^{+}$-expansion of $(\mathcal{K},\leq)$.

### v2

@@ 1. Introduction
The paper [12] of Kechris, Pestov and Todorčević established links between topological dynamics and structural Ramsey theory, with further developments in [15], [17], [2] (among others). We assume the reader is familiar with the background here, and briefly recall three key results, which we formulate for strong classes (classes of structures where we restrict to a particular subclass of permitted embeddings – see [5, Definition 2.1]):
@@
Theorem ([12, Theorem 10.8], [15, Theorem 5], [17, Theorem 5.7]).
@@
The paper [5], which was the starting point for the current paper, showed that classes of sparse graphs used in Hrushovski constructions ([9], [10]) demonstrate different behaviour to classes previously studied in the KPT context. A graph $A$ is $k$-sparse if for all finite $B\subseteq A$, we have $|E(B)|\leq k|B|$. It is well-known ([14], [5, Theorem 3.4]) that a graph is $k$-sparse iff it is $k$-orientable. We take $k=2$ for presentational simplicity.
We briefly describe the classes $\mathcal{C}_{0},\mathcal{C}_{1},\mathcal{C}_{F}$ of sparse graphs found in [4], [5]. Let $\mathcal{C}_{0}$ denote the class of finite $2$-sparse graphs. For $A,B\in\mathcal{C}_{0}$, we write $A\leq_{s}B$ if there exists a $2$-orientation of $B$ in which $A$ is successor-closed (by [7, Lemma 1.5], this is equivalent to another phrasing in terms of predimension). With this notion of $\leq_{s}$-substructure, we have the free amalgamation class $(\mathcal{C}_{0},\leq_{s})$ with Fraïssé limit $M_{0}$ (this structure, an “ab initio Hrushovski construction”, was first described in [10]).
We also have a “simplified” version of $M_{0}$, denoted by $M_{1}$ and first studied in [7]. This is the key structure we consider in this paper. Let $\mathcal{D}_{1}$ denote the class of finite $2$-oriented graphs with no directed cycles, and let $\mathcal{C}_{1}$ be the class of graph reducts of structures in $\mathcal{D}_{1}$. For $A,B\in\mathcal{C}_{1}$, write $A\leq_{1}B$ if there exists an expansion $B^{+}\in\mathcal{D}_{1}$ in which $A$ is successor-closed. Then $(\mathcal{C}_{1},\leq_{1})$ is again a free amalgamation class, and we write $M_{1}$ for its Fraïssé limit.
@@
Equivalently, by [12, Theorem 4.8], we have that $M$ has no coprecompact Ramsey expansion.
@@
Note that if $M(G)$ has a comeagre orbit, then so does any minimal $G$-flow (see [1]), so the above result shows again that $\operatorname{Aut}(M)$ for $M=M_{1},M_{0},M_{F}$ has non-metrisable universal minimal flow, using [2, Theorem 1.2]. In the context of the above result, T. Tsankov asked the following ([5, concluding remarks]):
@@
David Evans suggested that the author investigate the $\operatorname{Aut}(M)$-flow $\mathcal{LO}(M)$ of linear orders on $M$. We obtain the following result (the main result of this paper), for $M_{1}$, the “simplified version” of $M_{0}$:
@@
This result demonstrates that the phenomenon seen in [5, Theorem 5.2] occurs more generally for other flows on $M_{1}$, partially answering the question of Tsankov. We also note in passing that $M_{1}$ is $\omega$-saturated and its theory is $\omega$-stable (see [7]).
To prove 4.1, we take the class of finite ordered graphs which $\leq$-embed into some element of the minimal flow $Y$, and show that this class fails to have the weak amalgamation property – this gives the result, using 2.3. To show failure of the weak amalgamation property, we will use the Ramsey expansion given by the admissible orders, from [6, Section 3.1]. We discuss these in Section 3.
@@
Recall that we consider $M_{1}$ to be a “simplified version” of $M_{0}$. It would be interesting to know if there is an analogous “simplified version” of $M_{F}$ – if so, it may be possible to prove 4.1 for an $\omega$-categorical structure. See [16, Chapter 7].
@@ 2. Background
We briefly summarise the material required for Section 3 and 4.1. We assume thorough familiarity with [5] and the background and notation provided therein. This section contains no new material. (A reader looking for a less streamlined presentation may consult Chapter 1 of the author’s PhD thesis [16], again mostly based on [5].)
@@
2.1. Graphs and oriented graphs: notation
We write $E_{A}\subseteq A^{2}$ for the (symmetrised) edge set of a graph $A$ (where $(x,y)\in E_{A}$ iff $\{x,y\}$ is an undirected edge of $A$), and write $\rho_{A}\subseteq A^{2}$ for an orientation of $E_{A}$ (see [5, Definition 3.3]). A subgraph will be a first-order substructure, i.e. an induced subgraph. For an oriented graph $(A,\rho_{A})$, if $(x,y)\in\rho_{A}$, we write $xy\in\rho_{A}$, and for $x\in A$ we write $\operatorname{d_{{}_{+}}\!}(x)$ for the out-degree of $x$.
@@ 2.2. Sparse graphs: $\mathcal{C}_{0}$ and $\mathcal{C}_{1}$
Recall ([5, Definition 3.12]) that $\mathcal{C}_{0}$, $\mathcal{D}_{0}$ denote the classes of finite $2$-sparse graphs and finite $2$-oriented graphs, and that $\mathcal{C}_{0}$ is the class of graph reducts of $\mathcal{D}_{0}$. Let $\mathcal{D}_{1}$ be the class of finite $2$-oriented graphs with no directed cycles. By a slight abuse of terminology, we call a $2$-oriented graph with no directed cycles an acyclic $2$-oriented graph. Let $\mathcal{C}_{1}$ be the class of graph reducts of $\mathcal{D}_{1}$. We may consider $\mathcal{D}_{1}$, $\mathcal{C}_{1}$ as “simplified versions” of $\mathcal{D}_{0}$, $\mathcal{D}_{1}$: these classes are the key examples we are concerned with in this paper. They are originally from [7] and found in early preprint versions ([4, Definition 3.16]) of [5], though they do not appear in the published version.
@@
We may also define $\mathcal{C}_{1}$ directly: the class $\mathcal{C}_{1}$ consists of the finite graphs $A$ where every non-empty subgraph $B\subseteq A$ has a vertex of degree $\leq 2$. This follows from the fact ([7, Lemma 1.3]) that a finite graph $A$ has an acyclic $k$-orientation iff every non-empty subgraph $B$ has a vertex of degree $\leq k$ in $B$.
@@ 2.3. Sparse graphs: $\sqsubseteq_{s}$ and $\leq_{1}$
We now describe the distinguished notions of embedding used to define the particular strong classes $(\mathcal{C}_{1},\leq_{1})$, $(\mathcal{D}_{1},\sqsubseteq_{s})$. (Śee [5, Definition 2.1] for strong classes and [5, Section 3.4] for basic lemmas regarding $\sqsubseteq_{s}$. The case of $\leq_{1}$ is originally from [8].)
@@
For an oriented graph $A$ and $B\subseteq A$, we write $B\sqsubseteq_{s}A$ to mean that $B$ is successor-closed in $A$. For $B\subseteq A$, the successor-closure $\operatorname{scl}(B)$ is the smallest successor-closed subset of $A$ containing $B$.
@@
Let $A,B\in\mathcal{C}_{1}$ with $A\subseteq B$. We write $A\leq_{1}B$ if there exists an acyclic $2$-orientation $B^{+}\in\mathcal{D}_{1}$ of $B$ in which the induced orientation $A^{+}\in\mathcal{D}_{1}$ on $A$ has $A^{+}\sqsubseteq_{s}B^{+}$. By an argument entirely analogous to [5, Section 3.4 and Lemma 4.8], it is easy to show the following.
Fact ([4, Theorem 3.17]).
@@
$(\mathcal{C}_{1},\leq_{1})$ and $(\mathcal{D}_{1},\sqsubseteq_{s})$ are strong classes with free amalgamation, and the class $(\mathcal{D}_{1},\sqsubseteq_{s})$ is both a strong and a reasonable expansion of $(\mathcal{C}_{1},\leq_{1})$.
(Recall the definition of a strong expansion and a reasonable expansion from [5, Definition 2.9, Definition 2.14].)
Let $M_{1}$ be the Fraïssé limit of $(\mathcal{C}_{1},\leq_{1})$ and let $G_{1}=\operatorname{Aut}(M_{1})$. We now discuss the technical framework used to analyse minimal subflows of $\mathcal{LO}(M_{1})$, the $G_{1}$-flow of linear orders on $M_{1}$.
2.4. The order expansion of an amalgamation class
Let $(\mathcal{K},\leq)$ be an amalgamation class of $L$-structures, and let $L^{\prec}=L\cup\{\prec\}$ with $\prec$ binary. Let $\mathcal{K}^{\prec}$ be the class of $L^{+}$-structures $(A,\prec_{A})$, where $A\in\mathcal{K}$ and $\prec_{A}$ is a linear order on $A$. For $(A,\prec_{A}),(B,\prec_{B})\in\mathcal{K}^{\prec}$, write $(A,\prec_{A})\leq(B,\prec_{B})$ if $A\leq B$ and $\prec_{A}$ is the restriction of $\prec_{B}$ to $A$. We call $(\mathcal{K}^{\prec},\leq)$ the order expansion of $(\mathcal{K},\leq)$. It is straightforward to check that $(\mathcal{K}^{\prec},\leq)$ is a strong class which is both a strong and reasonable expansion of $(\mathcal{K},\leq)$.
2.5. Flows from reasonable expansions